# Towards Reliable Benchmarking: A Contamination Free, Controllable Evaluation Framework for Multi-step LLM Function Calling

### v1

@@ Abstract
As language models gain access to external tools through structured function calls, they become increasingly more capable of solving complex, multi-step tasks. However, existing benchmarks for tool-augmented language models (TaLMs) provide insufficient control over factors such as the number of functions accessible, task complexity, and input size, and remain vulnerable to data contamination. We present FuncBenchGen, a unified, contamination-free framework that evaluates TaLMs by generating synthetic multi-step tool-use tasks to stress-test TaLMs. The key idea is to cast tool use as traversal over a hidden function-dependency DAG where nodes are function calls and an edge between nodes represents one function consuming the output of another. Given a set of external function schemas, initial variable values, and a target variable, models must compose the correct call sequence to compute the target variable. FuncBenchGen allows users to precisely control task difficulty (e.g., graph size, dependency depth, and distractor functions) while avoiding pretraining/test-time leakage.
We apply our FuncBenchGen framework to evaluate seven open and closed LLMs on tool use tasks of varying difficulty. Reasoning-optimized models consistently outperform general-purpose models with GPT-5 significantly outperforming other available models. Performance declines sharply as dependency depth increases. Furthermore, connected distractors—irrelevant functions sharing type-compatible variables with relevant functions—prove especially difficult to handle.
We find that strong models often make syntactically valid function calls but propagate incorrect or stale argument values across steps, revealing brittle state tracking by LLMs in multi-turn tool use. Motivated by this observation, we introduce a simple mitigation strategy that explicitly restates prior variable values to the agent at each step. Surprisingly, this lightweight change yields substantial gains across models. e.g., yielding an improvement in success rate from 62.5% to 81.3% for GPT-5, without modifying the underlying architectures or training.
@@ 2 Related Work
Recent LLMs demonstrate impressive multi-step reasoning capabilities across various domains, including mathematical problem solving (Davoodi et al., 2025), multi-document reasoning (Maekawa et al., 2025), and commonsense reasoning (Yu et al., 2025). Despite these advances, the robustness of their multi-step reasoning capabilities remains unclear when models are required to interact with external tools.
@@
Models such as ToolLLM (Qin et al., 2024), Gorilla (Patil et al., 2024), and ToolACE (Liu et al., 2025a) highlight the importance of equipping LLMs with access to vast API collections to tackle real-world use cases. Tool-Planner (Liu et al., 2025b) and ToolDial (Shim et al., 2025) further extend these capabilities by incorporating planning and multi-turn dialogue for more sophisticated tool-use by using LLMs’ multi-step reasoning capabilities. Meanwhile, LongFuncEval (Kate et al., 2025) and ComplexFuncBench (Zhong et al., 2025) focus on evaluating LLMs’ ability to handle long-context and multi-step function calling scenarios. Despite these advances, the evaluation datasets used often lack generality and controllability, limiting their utility for systematic analysis.
@@
Robustness and contamination issues in benchmark construction pose significant challenges. Datasets such as (Mirzadeh et al., 2025; Shojaee et al., 2025) have addressed the limitations of LLM reasoning evaluation due to dataset contamination or task leakage, e.g., when the LLM training set includes a specific tool or task included in the test set. Han et al. (2025) have shown data contamination can happen during web search tool-use. Though these studies emphasize the importance of creating bias-free and contamination-free tasks, they do not specifically address the challenges in multi-step function calling scenarios. In contrast, our work seeks to bridge this gap by offering a principled, automated framework for function set and task generation, supporting controlled stress-testing and robust analysis of TaLM capabilities.
@@ 3.2 Benchmark Generation Process
The framework takes a two-stage approach: 1) core node creation and 2) irrelevant node addition. First, we create a node sequence of length $d$ to ensure that the generated graph contains a valid path from input variables to the target variable with the required number of function calls. Then, we iteratively add the remaining $n^{\text{core}}-(d+1)$ core nodes by randomly adding new parent nodes to the existing core nodes, while ensuring acyclicity and dependency depth. Second, we add irrelevant nodes according to the specified connection type. For connected irrelevant nodes (CINs), we randomly select nodes from the existing core nodes and add the irrelevant nodes as their children. For disconnected irrelevant nodes (DINs), we create isolated nodes and then add at most $\lfloor n^{\text{dis}}/2\rfloor$ edges between them while ensuring acyclicity, to simulate the possible function dependency between irrelevant nodes.
@@
Functions are linked through semantic type and sub-type matching rather than variable names, serving as a lightweight proxy for the semantic reasoning that connects functions in real-world scenarios. In practice, if a function’s input type and sub-type matches another’s output type and sub-type, the two are connected. See A.2 for details.
@@ 3.3 Real-world Applicability
Many real-world function calling scenarios can be mapped directly to DAG structures. For example, a user may ask an LLM agent to find a sightseeing tour near their hotel tomorrow and reserve a taxi for them an hour after the tour ends (example from ComplexFuncBench (Zhong et al., 2025)). Here, an agent must: 1) find the hotel’s location, 2) use the location to find nearby tours, 3) filter nearby tours by date and availability, and 4) reserve a taxi at the correct time and place. This workflow naturally forms a dependency graph where each function depends on specific outputs from previous calls—tours cannot be searched for without the hotel’s location, a tour cannot be selected without the tour search results, etc. Our framework abstracts this complexity into a controllable evaluation setting while preserving the essential challenge: models must infer these dependencies from function signatures alone, just as they would when interacting with unfamiliar APIs in deployment.
@@ 3.4 Implementation Details
We implement an interaction protocol between the LLM agent and the function execution environment. At each step, the model can make multiple function calls, then receive their outputs, and continue until it makes no further calls. An example of the model input is shown in Appendix A.3.
@@ 4.1 Setup
We evaluate various LLMs with reasoning-optimized and general-purpose capabilities. Reasoning models include both closed and open models: GPT-5, GPT-5-mini, Gemini-2.5-Pro, Gemini-2.5-Flash, Qwen3-235B22A (Qwen3 for short). General models include GPT-4.1 and GPT-4.1-mini. The model details are summarized in Appendix A.1. We set the thinking budgets of all the reasoning models to their default, e.g., medium for GPT-5, and set the temperature and top-p parameters to $0.0$ and $1.0$, respectively, for all our experiments. [^1] Since GPT-5 and GPT-5-mini do not support the temperature parameter, we use their default settings.
@@ 4.2 Main Results (RQ1)
Success
Rate
ACs
ACs
Success
Rate
ACs
ACs
Success
Rate
ACs
ACs
@@ 4.3 Effect of Irrelevant Functions (RQ2)
Figure 3 shows that CINs have the most negative impact on model performance compared to other types across most tested LLMs. We speculate that this is because CINs share variables with core nodes, making it challenging for models to distinguish relevant function paths from irrelevant ones.
@@ 4.4 Effect of Function Dependency Depth (RQ3)
We break down the results by the dependency depth in Figure 4. We also show the results with $95\%$ confidence intervals and more numbers of core nodes in Appendix B.
@@ 4.6 Failure Analysis and Mitigation Strategy (RQ5)
In analyzing model failures in function calling, we use complete execution traces from our in-house executor, which logs all calls, arguments, returns, and state updates, enabling deterministic, annotation-free attribution. The executor enforces four sequential, machine-checked predicates—name resolution, schema conformance, dataflow availability, and value consistency—designed to be minimal and collectively exhaustive. We categorize the failure cases into four types: 1) Function Not Found: The model attempts to call a function that does not exist in the provided function list. 2) Wrong Number of Inputs: The model provides more or fewer input arguments than the function schema allows. 3) Value Not Yet Known: The model tries to use a variable whose value has not been established through prior function calls or initial inputs. 4) Incorrect Value: The model uses a variable with an incorrect value, which has been established through prior function calls or initial inputs.
@@ 5 Conclusion
We present FuncBenchGen, a novel framework for generating multi-step function calling benchmarks that address key limitations in existing evaluations. By enabling controllable complexity and contamination-free tasks, FuncBenchGen provides a robust platform for systematically assessing LLMs’ function calling capabilities. Our extensive experiments reveal significant performance gaps between reasoning-optimized and general-purpose models. Results also highlight challenges posed by irrelevant functions and long call sequences – even GPT-5 struggles with longer function call sequences. We identify common failure types, and based on this analysis propose a simple yet effective mitigation strategy that significantly improves performance across various models. Overall, FuncBenchGen offers a valuable tool for advancing research into LLM function calling capabilities.
@@
We have made several efforts to ensure the reproducibility of our work. First, the design principles, methodology, and evaluation setup for FuncBenchGen are detailed in Section 3. All models used in the experiments are well-documented in Appendix A.1, with the prompt example in Appendix A.3. Upon paper acceptance, we will release the framework, codebase, and detailed documentation for both the experimental setup and data generation to promote ease of replication.
@@
Appendix A Additional Experiment Details
A.1 Model Details
@@ A.2 Types and Subtypes
A.3 Example Model Input
@@
Appendix B Additional Results
B.1 Results of No Extra Setting
@@
Success
Rate
ACs
ACs
Success
Rate
ACs
ACs
Success
Rate
ACs
ACs
@@
B.2 Results of Dependency Depth with Error Bars
@@
B.3 Results of Other Dependency Depths

### v2

@@ Abstract
Existing benchmarks for tool-augmented language models (TaLMs) lack fine-grained control over task difficulty and remain vulnerable to data contamination. We present FuncBenchGen, a unified, contamination-free framework that evaluates TaLMs by generating synthetic multi-step tool-use tasks to stress-test TaLMs. The key idea is to cast tool use as traversal over a hidden function-dependency DAG where models must infer the correct sequence of calls to compute a target value. FuncBenchGen allows precise control over task difficulty (e.g., graph size, dependency depth, and distractor functions) while avoiding pretraining/test-time leakage. Our evaluation demonstrates reasoning-optimized models consistently outperform general-purpose models with GPT-5 significantly outperforming other available models. Performance declines sharply as dependency depth increases. Furthermore, connected distractors—irrelevant functions sharing type-compatible variables with relevant functions—prove especially difficult to handle. Also, strong models often make syntactically valid function calls but propagate incorrect or stale argument values across steps, revealing brittle state tracking by LLMs in multi-turn tool use. Motivated by this observation, we introduce a simple mitigation strategy that explicitly restates prior variable values to the agent at each step. Surprisingly, this lightweight change yields substantial gains across models. e.g., yielding an improvement in success rate from 62.5% to 81.3% for GPT-5.
@@ 2 Related Work
Recent LLMs demonstrate impressive multi-step reasoning capabilities across various domains, including mathematical problem solving (Davoodi et al., 2025), multi-document reasoning (Maekawa et al., 2025), and commonsense reasoning (Yu et al., 2025). Other work has focused on the rigor and logical validity of this reasoning. For example, FOLIO (Han et al., 2024) and Multi-LogiEval (Patel et al., 2024) evaluate complex multi-step logical reasoning using first-order logic. Despite these advances, the robustness of their multi-step reasoning capabilities remains unclear when models are required to interact with external tools.
Code Generation and Reasoning
The ability to use tools is fundamentally linked to a model’s ability to reason about code execution. This foundational capability has been evaluated by recent benchmarks (Gu et al., 2024; Chen et al., 2025). While these benchmarks focus on reasoning about code, our work evaluates the applied task of using code as black-box external tools.
@@
Models such as ToolLLM (Qin et al., 2024), Gorilla (Patil et al., 2024), and ToolACE (Liu et al., 2025a) highlight the importance of equipping LLMs with access to vast API collections to tackle real-world use cases. Tool-Planner (Liu et al., 2025b), ToolDial (Shim et al., 2025), HammerBench (Wang et al., 2025) further extend these capabilities by incorporating planning and multi-turn dialogue for more sophisticated tool-use by using LLMs’ multi-step reasoning capabilities. Meanwhile, LongFuncEval (Kate et al., 2025) and ComplexFuncBench (Zhong et al., 2025) focus on evaluating LLMs’ ability to handle long-context and multi-step function calling scenarios. Despite these advances, the evaluation datasets used often lack generality and controllability, limiting their utility for systematic analysis.
@@
Robustness and contamination issues in benchmark construction pose significant challenges. Datasets such as (Mirzadeh et al., 2025; Shojaee et al., 2025) have addressed the limitations of LLM reasoning evaluation due to dataset contamination or task leakage, e.g., when the LLM training set includes a specific tool or task included in the test set. Han et al. (2025) have shown data contamination can happen during web search tool-use. White et al. (2025) have introduced LiveBench, a contamination-limited benchmark that is refreshed regularly and spans multiple domains such as math, reasoning, and instruction following tasks. Though these studies emphasize the importance of creating bias-free and contamination-free tasks, they do not specifically address the challenges in multi-step function calling scenarios. In contrast, our work seeks to bridge this gap by offering a principled, automated framework for function set and task generation, supporting controlled stress-testing and robust analysis of TaLM capabilities.
@@ 3.2 Benchmark Generation Process
The framework takes a two-stage approach: 1) core node creation and 2) irrelevant node addition. First, we create a node sequence of length $d$ to ensure that the generated graph contains a valid path from input variables to the target variable with the required number of function calls. Then, we iteratively add the remaining $n^{\text{core}}-(d+1)$ core nodes by randomly adding new parent nodes to the existing core nodes, while ensuring acyclicity and dependency depth. Second, we add irrelevant nodes according to the specified connection type. For connected irrelevant nodes (CINs), we randomly select nodes from the existing core nodes and add the irrelevant nodes as their children. For disconnected irrelevant nodes (DINs), we create isolated nodes and then add at most $\lfloor n^{\text{dis}}/2\rfloor$ edges between them while ensuring acyclicity, to simulate the possible function dependency between irrelevant nodes. We provide illustrative examples of the generated graphs and more graph generation details in Appendix A.1.
@@
Functions are linked through semantic type and sub-type matching rather than variable names, serving as a lightweight proxy for the semantic reasoning that connects functions in real-world scenarios. In practice, if a function’s input type and sub-type matches another’s output type and sub-type, the two are connected. See B.2 for details.
@@ 3.3 Real-world Applicability
Figure 3: Graphical representation of a real tool-calling example sequence. Many tool-calling applications can be represented as DAGs in this way.
Many real-world function calling scenarios can be mapped directly to DAG structures. For example, a user may ask an LLM agent to find a sightseeing tour near their hotel tomorrow and reserve a taxi for them an hour after the tour ends (example from ComplexFuncBench (Zhong et al., 2025)). Here, an agent must: 1) find the hotel’s location, 2) use the location to find nearby tours, 3) filter nearby tours by date and availability, and 4) reserve a taxi at the correct time and place. Figure 3 illustrates the function dependency graph for this example. This workflow naturally forms a dependency graph where each function depends on specific outputs from previous calls—tours cannot be searched for without the hotel’s location, a tour cannot be selected without the tour search results, etc. Our framework abstracts this complexity into a controllable evaluation setting while preserving the essential challenge: models must infer these dependencies from function signatures alone, just as they would when interacting with unfamiliar APIs in deployment.
@@ 3.4 Implementation Details
We implement an interaction protocol between the LLM agent and the function execution environment. At each step, the model can make multiple function calls, then receive their outputs, and continue until it makes no further calls. An example of the model input is shown in Appendix B.3.
@@ 4.1 Setup
We evaluate various LLMs with reasoning-optimized and general-purpose capabilities. Reasoning models include both closed and open models: GPT-5, GPT-5-mini, Gemini-2.5-Pro, Gemini-2.5-Flash, Qwen3-235B22A (Qwen3 for short). General models include GPT-4.1 and GPT-4.1-mini. The model details are summarized in Appendix B.1. We set the thinking budgets of all the reasoning models to their default, e.g., medium for GPT-5, and set the temperature and top-p parameters to $0.0$ and $1.0$, respectively, for all our experiments. [^1] Since GPT-5 and GPT-5-mini do not support the temperature parameter, we use their default settings.
@@ 4.2 Main Results (RQ1)
Success Rate ACs (Succ.) ACs (Fail.) Success Rate ACs (Succ.) ACs (Fail.) Success Rate ACs (Succ.) ACs (Fail.)
@@ 4.3 Effect of Irrelevant Functions (RQ2)
Figure 4 shows that CINs have the most negative impact on model performance compared to other types across most tested LLMs. We speculate that this is because CINs share variables with core nodes, making it challenging for models to distinguish relevant function paths from irrelevant ones. The similar trend of the performance drop is observed in other reasoning tasks (Shi et al., 2023) when distracting information is present.
@@ 4.4 Effect of Function Dependency Depth (RQ3)
We break down the results by the dependency depth in Figure 5. We also show the results with $95\%$ confidence intervals and more numbers of core nodes in Appendix C.
@@ 4.6 Failure Analysis and Mitigation Strategy (RQ5)
In analyzing model failures in function calling, we use complete execution traces from our in-house executor, which logs all calls, arguments, returns, and state updates, enabling deterministic, annotation-free attribution. The executor enforces four sequential, machine-checked predicates—name resolution, schema conformance, dataflow availability, and value consistency—designed to be minimal and collectively exhaustive. We categorize the failure cases into four types: 1) Function Not Found: The model attempts to call a function that does not exist in the provided function list. 2) Wrong Number of Inputs: The model provides more or fewer input arguments than the function schema allows. 3) Value Not Yet Known: The model tries to use a variable whose value has not been established through prior function calls or initial inputs. 4) Incorrect Value: The model uses a variable with an incorrect value, which has been established through prior function calls or initial inputs. These failure types cover all systematically detectable errors, as they correspond to the hierarchical prerequisites required for any valid function execution: 1) function existence errors, 2) argument mismatch errors, 3) incorrect call sequence errors, and 4) state-tracking errors. These failure types can be considered in sequence (1–4) using if-branches, and any error arising from a function call will fall into one of these categories.
@@ 4.6.1 Mitigation Strategy
4.7 Case Study of Function Calling Failure
To illustrate how LLMs fail in function calling tasks, we present a case study of GPT-5’s failure examples in Figure 9. In Figure 9(a), GPT-5 incorrectly calls irrelevant functions such as func_pbb and func_ped in the seventh and eighth calls, respectively. Then, we observe that GPT-5 invokes the core function func_nss with wrong input values in the ninth call, i.e., it uses correct variable values from func_puf and func_eni but also uses an incorrect value from func_pbb instead of func_hoj even though it was successfully called in the fifth call. In Figure 9(b), GPT-5 misses required function calls such as func_rgy and func_sqo (the blue nodes without a number at the top-right) throughout the entire function calling sequence. As a result, GPT-5 attempts to call the target function func_xrx with irrelevant values from func_qcu and func_otq in the fifth and ninth call. These examples highlight the challenges LLMs face in accurately navigating function call sequences, particularly in avoiding irrelevant functions and ensuring correct input values for core functions. We provide a full function calling sequence of the example in Appendix C.4.
(a) A failure example in which GPT-5 calls irrelevant functions and invoke core functions with wrong input values.
(b) A failure example in which GPT-5 misses required function calls and attempt to call the target function with irrelevant values.
Figure 9: Analysis of GPT-5 function calling errors. The generated graphs feature a depth of 5, comprising 10 core nodes, 10 CINs, and 0 DINs across various random seeds. The number at the top-right of each node denotes the sequential order of function calls made by GPT-5, e.g., "2" indicates the second function called. The color of the number indicates the result: black for correct calls, green for irrelevant nodes (no error), and red for errors.
@@ 5 Conclusion
We present FuncBenchGen, a novel framework for generating multi-step function calling benchmarks that address key limitations in existing evaluations. By enabling controllable complexity and contamination-free tasks, FuncBenchGen provides a robust platform for systematically assessing LLMs’ function calling capabilities. Our extensive experiments reveal significant performance gaps between reasoning-optimized and general-purpose models. Results also highlight challenges posed by irrelevant functions and long call sequences – even GPT-5 struggles with longer function call sequences. We identify common failure types, and based on this analysis propose a simple yet effective mitigation strategy that significantly improves performance across various models. Overall, FuncBenchGen offers a valuable tool for advancing research into LLM function calling capabilities. Incorporating more complex control flows such as conditional logic and iteration would be our interesting future direction.
@@
We have made several efforts to ensure the reproducibility of our work. First, the design principles, methodology, and evaluation setup for FuncBenchGen are detailed in Section 3. All models used in the experiments are well-documented in Appendix B.1, with the prompt example in Appendix B.3.
@@
Appendix A Additional Details of Graph Generation
A.1 Graph Generation Process Detail
Figure 10: Graph generation process.
We illustrate the graph generation process in Figure 10. As described in Section 3, the graph generation consists of two main steps: core node creation and irrelevant node addition. In the core node creation step, we first generate a initial graph that is a sequence of nodes based on the specified number of dependency depth as shown in the left in the figure. Then, we iteratively add remaining nodes to the graph until reaching the specified number of core nodes while ensuring that the maximum dependency depth is not exceeded (see the middle in the figure). To increase the diversity, we randomly add up to $2n^{\text{core}}$ edges between core nodes while maintaining the acyclic property of the graph. If adding the new node would exceed the maximum dependency depth, we discard that choice and try again. In the irrelevant node addition step, we add irrelevant nodes based on the specified connection type (see the right in the figure). For connected irrelevant nodes, we randomly select parent nodes from the existing core nodes and link the new irrelevant nodes to it as their children. For disconnected irrelevant nodes, we simply add the new irrelevant node without linking it to any core nodes.
A.2 Example Generated Graphs
(a) Seed=0.
(b) Seed=$100$.
Figure 11: Example DAGs with depth 5, 10 core nodes, 10 CINs, 0 DINs, and different random seeds.
To provide concrete examples of our generated graphs, we present figures illustrating DAGs with varying depths and numbers of core and irrelevant nodes used in our experiments. Specifically, Figure 11 displays generated DAGs with a depth of 5, 10 core nodes, 10 CINs, and 0 DINs, generated using different random seeds (0 and 100 in Figures 11(a) and 11(b), respectively). These examples demonstrate how the overall structures of the DAGs can vary significantly—for instance, one graph features two first-layer core nodes while the other has four—while maintaining the same depth and total number of nodes. In our experiments, we use five different random seeds and aggregate the results to ensure robustness against these structural variations.
(a) Depth=3.
(b) Depth=11.
Figure 12: Example DAGs with different depths, 20 core nodes, 5 CINs, and 5 DINs.
Next, we provide generated DAGs with depth 3 and 11, 20 core nodes, 5 CINs, and 5 DINs in Figure 12, to demonstrate how the structure changes with varying depths while keeping the number of core and irrelevant nodes constant. Figures 12(a) and 12(b) illustrate these graphs, respectively. The depth 3 graph exhibits a wider structure, while the depth 11 graph shows a longer chain of dependencies. These examples highlight the flexibility of our graph generation process in creating diverse structures that can effectively test the function calling capabilities of LLMs under different complexity levels.
Appendix B Additional Experiment Details
B.1 Model Details
@@ B.2 Types and Subtypes
B.3 Example Model Input
@@
Appendix C Additional Results
C.1 Results of No Extra Setting
@@
Success Rate ACs (Succ.) ACs (Fail.) Success Rate ACs (Succ.) ACs (Fail.) Success Rate ACs (Succ.) ACs (Fail.)
@@
C.2 Results of Dependency Depth with Error Bars
@@
C.3 Results of Other Dependency Depths
@@
Figure 16 shows the comparison between the baseline and the proposed mitigation strategy with $10$ core nodes. We observe similar trends to those in Figure 8.
@@
Figure 16: Comparison between the baseline and the proposed mitigation strategy. The number of core nodes is set to $10$. The results are averaged across $\{0,10,20,40\}$ CINs, with 5 trials each.
C.4 Function Calling Sequence Example
Figure 17 shows a full function calling sequence example of the case study in Figure 9(a).
Figure 17: Detailed function calling sequence. The left block shows the input to the function call which the model generates, and the right block shows the output from the function call. Core nodes are highlighted in blue, connected irrelevant nodes (CINs) in green, and failed function calls in red. Errors are annotated below the output values.